A chemist is making a solution that is to be 55% chlorine. He has one solution that is 40% chlorine and another that is 65% chlorine. He wants 100 liters of the final solution. How much of each should he mix?

Answer by checkley75(3666)

.4x+.65(100-x)=100*.55
.4x+65-.65x=55
-.25x=55-65
-.25x=-10
x=10/.25
x=40 liters of 40% chlorine solution
100-40=60 liters of 65% solution.
proof
.4*40+.65*60=55
16+39=55
55=55